ETAPS’99 Report

Y. Annie Liu and Scott D. Stoller Computer Science Department, Indiana University

April 1999

The European Joint Conferences on Theory and Practice of Software (ETAPS) is a loose and open confederation of existing and new conferences and other events that has become the primary European forum for academic and industrial researchers working on topics relating to Software Science. By bringing these conferences and events together, ETAPS fosters interaction among researchers, promotes cross-fertilization of ideas from different technical areas, and reduces travel overhead. ETAPS was founded in 1998 after three years of discussions and three more years of active planning and is now an annual meeting that takes place in Europe each spring. After the first ETAPS in Lisbon in 1998, ETAPS’99 was held in Amsterdam in 1999, March 20th through 28th.

ETAPS’99 comprised 5 conferences, 5 satellite events, 7 invited lectures, 2 invited tutorials, and 4 contributed tutorials. The conferences were:

• CC’99, International Conference on Compiler Construction, Stefan Jähnichen, chair

• ESOP’99, European Symposium on Programming, Doaitse Swierstra, chair

• TACAS’99, Tools and Algorithms for the Construction and Analysis of Systems, Rance Cleaveland, chair

• FASE’99, Fundamental Approaches to Software Engineering, Jean-Pierre Finance, chair

• FoSSaCS’99, Foundations of Software Science and Computation Structures, Wolfgang Thomas, chair

The conferences contained more than 100 technical papers. Each conference had a separate program committee and a separate proceedings. All of the proceedings were published by Spring-Verlag in the LNCS series, as volumes 1575-1579. One conference fee entitled a participant to attend all of the conferences and choose any two of the proceedings.

Satellite events included:

• CMCS’99, International Workshop on Coalgebraic Methods in Computer Science

• AS’99, International Workshop on Action Semantics

• CWI Soiree, Invited Speakers: Bjarne Stroustrup, Edmund Clarke

• WAGA’99, International Workshop on Attribute Grammars and Their Applications

• CoFI, Common Framework Initiative

The workshops offered dozens of talks. The CWI Soiree attracted a full house, even though it happened late in the evening; the two invited talks were What is an object and what isn’t? (Stroustrup) and Model checking: past, present, and future (Clarke).

The conferences were bound together by a common program of invited lectures and tutorials, covering a wide range of topics of broad interest. The invited speakers were:

• Gilad Bracha (Sun Microsystems, USA), Making Java easier to type, and easier to type

• Paul Hudak (Yale University, USA), Functional reactive programming

• Herbert Weber (Technische Universität Berlin, Germany), Continuous engineering of information and communication infrastructure

• Martín Abadi (Compaq Systems Research Center, Palo Alto, USA), Security protocols and specifications

• Jeff Kramer (Imperial College, UK), Modelling for mere mortals

• Jean-Claude Laprie (LAAS Toulouse, France), Software-based critical systems

• Javier Esparza (Technische Universität Mnchen, Germany), Grammars as processes

These invited lectures packed the large lecture hall early in the mornings.

Another event that unified the meeting was a panel discussion entitled Supporting software research: what does PITAC mean to you?. Susan Graham (UC Berkeley), a PITAC member, led off with an overview of the context and conclusions of the PITAC report. The discussion brought out the differences and similarities of the software research environments in Europe and the United States.

The succinct opening of ETAPS’99 by Don Sannella, chair of the ETAPS’98, ETAPS’99, and ETAPS’2000 steering committees, conveyed an important message, that is, by bringing together a variety of conferences and related events, ETAPS encourages theory with a practical motivation on one hand and soundly-based practice on the other. Individual interests drove many of the participants to move among the parallel sessions of the conferences.

ETAPS’99 was organized locally by Jan Bergstra of CWI and the University of Amsterdam together with Frans Snijders of CWI. The conferences were held at the Koninklijk Instituut voor de Tropen (Royal Tropical Institute), conveniently located a short tram ride away from the city center. The main building dates back to the 1920’s. The meeting rooms are richly decorated, with statues perched high among the rafters in one room, and a stone Buddha sitting placidly (and safely) near floor level in another. The participants could also enjoy the Institute’s Tropical Museum.

The first day of conferences wound down with an evening reception. The Mayor and Aldermen of the City of Amsterdam welcomed ETAPS’99 by inviting the attendees to a reception in the City Hall. After three busy days of conferences, everyone was ready to enjoy an evening aboard the Kapitein Kok, a traditional Dutch wheelboat touring the IJsselmeer (formerly a part of the North Sea, now a lake separated from it by a dike). The winners of the best-paper awards were announced and congratulated after the dinner buffet and dessert buffet.

ETAPS’2000 will be in Berlin in March 25 through April 2, 2000, and will host the same five conferences along with tool demonstrations, tutorials, and a number of satellite events. Information about ETAPS’2000 can be found at http://iks.cs.tu-berlin.de/etaps2000/.

Annie Liu and Scott Stoller can be reached at {liu,stoller}@cs.indiana.edu.